OpenStudy (anonymous):

A world class sprinter can burst out of the blocks to essentially top speed (about 11.5m/s) in the first 15.0m of the race. What is the average acceleration of this sprinter, and how long does it take to reach that speed?

OpenStudy (anonymous):

Start with change in kinetic energy due to force (F) and distance. KE = (1/2) m v^2 = (F) (d) Get v^2 by itself, noting the F/m = a acceleration v^2 = 2 a d a = (v^2)/ 2 d reached top speed when went s = 15 m = (1/2) a t^2, so you can find t from this.